OpenStudy (anonymous):

Hey Guys! can you help me with this? Father can paint the house alone in 8 hours and it takes his son 12 hours to do the same job. How many hours does it take if they work together to complete the job?

OpenStudy (anonymous):

IDK how to do this, sorry.

OpenStudy (anonymous):

its ok @mipie13

OpenStudy (anonymous):

4.8 hours

OpenStudy (wolf1728):

There is a formula for this: f = 8 s = 12 time = 8*12 / 8+12 = 96 / 20 = 4.8 hours

OpenStudy (anonymous):

Father can do 1/8 of the job in 1 hours. His son can do 1/12 of the job in 1 hours. So in two hours, they can do 1/8+1/12 of the job = 5/24 of the job.They can do 5/24 of the job in 1 hour, times each side by 4.8 , you get 4.8 hours

OpenStudy (anonymous):

father in one hour do 1/8 of work and son do 1/12 of work.if both work in x hours :we have ;1/12(x)+1/8(x)=1 2/24(x)+3/24(x)=1 x=24/5=4.8 hours

OpenStudy (anonymous):

@fazeli where did the 24 came from?

OpenStudy (anonymous):

is it the LCM?

OpenStudy (anonymous):

\[\frac{ 1 }{ 12}x+\frac{ 1 }{ 8 }x=1\] \[\frac{ 2 }{ 24 }x+\frac{ 3 }{ 24}x=1\]

OpenStudy (anonymous):

4.8

OpenStudy (anonymous):

thanks @fazeli and @josjos also at @wolf1728